OpenStudy (anonymous):

Determine the number and type of solutions for x2 + 9x + 7 = 0

OpenStudy (anonymous):

can you determine the discriminant for me?

OpenStudy (anonymous):

o, idk what that is...

OpenStudy (anonymous):

in the equation \(\large ax^2+bx+c=0 \), the discriminant is \(\large D=b^2-4ac \) so in your equation \(\large x^2+9x+7=0 \), identify a, b, and c and plug it into that formula..:)

OpenStudy (anonymous):

what did you get for D= ?

OpenStudy (anonymous):

is it two rational?

OpenStudy (anonymous):

why do you say that?

OpenStudy (anonymous):

because it dosnt come up with any - and i thought negatives were irrational, and it isnt a 0 so i figured its rational!

OpenStudy (anonymous):

yep... you're correct

OpenStudy (anonymous):

yeah, it's positive so there are 2 real solutions

OpenStudy (anonymous):

r u sure, is it that or two complex??

OpenStudy (anonymous):

if D is negative, you have complex roots if D is positive, you have real roots if D=0 you have 1 real root